### v1

# Gravitational Waves from Type-I Strings in a Neutrino Mass Model

Abstract
In this work, we propose a novel realization of type-I cosmic strings arising from the spontaneous breaking of an extended gauge symmetry $SU(2)_{R}\times U(1)_{B-L}$ in the context of a low-scale split seesaw mechanism for neutrino mass generation. We demonstrate that the split seesaw framework, which explains the smallness of neutrino masses, naturally motivates a small scalar self-coupling $\lambda$. This intrinsically links the neutrino mass generation mechanism to the formation of type-I cosmic strings, where the gauge coupling dominates over the scalar self-coupling ($\beta\equiv\lambda/2g^{2}<1$). We explore the cosmological implications of these strings, including their gravitational wave signatures that are testable in current and future experiments. Our findings establish a compelling and testable connection between neutrino mass generation and cosmic string phenomenology in an underexplored region of parameter space.
I Introduction
Understanding the origin of neutrino masses and the possible remnants of high-scale symmetry breaking in the early universe are central challenges in modern particle physics and cosmology. Left-right (LR) symmetric models, based on the gauge group
$SU(3)_{C}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L},$
offer a compelling framework to address these questions. They naturally incorporate right-handed (RH) neutrinos as gauge singlets under the Standard Model, but charged under $SU(2)_{R}\times U(1)_{B-L}$, and thus enable the implementation of the seesaw mechanism for neutrino mass generation [1, 2]. Moreover, they provide a natural origin for parity violation in weak interactions, embedding it within a spontaneously broken symmetry [3].
Among the various implementations of the seesaw mechanism, the split seesaw scenario [4] provides an elegant explanation for the tiny active neutrino masses through geometric suppression in higher-dimensional theories. In this framework, RH neutrinos are localized at different positions in a compactified extra dimension. The overlap of their wavefunctions with the Standard Model brane results in exponentially suppressed Yukawa couplings, without requiring ultra-small fundamental parameters. This separation not only generates a realistic neutrino mass spectrum but also enables one of the RH neutrinos to be long-lived and a viable dark matter candidate [5, 6].
A critical prediction of LR symmetric theories is the spontaneous breaking of the gauge symmetry $SU(2)_{R}\times U(1)_{B-L}\rightarrow U(1)_{Y}$, which is typically achieved by the vacuum expectation value (VEV) of a RH triplet scalar $\Delta_{R}$. This symmetry breaking can lead to the formation of topological defects, specifically cosmic strings (CSs), if the vacuum manifold contains non-trivial first homotopy groups [7, 8]. In our model, this breaking gives rise to type-I CSs, characterized by a scalar self-coupling $\lambda$ much smaller than the gauge coupling squared $g^{2}$, i.e., $\beta\equiv\lambda/g^{2}\ll 1$ [9]. The formation of type-I CSs from supersymmetric flat directions has been explored in [10]. Such strings are attractive in nature and can influence early-universe dynamics and gravitational wave (GW) backgrounds.
Furthermore, the large VEV, $\nu_{R}$ associated with $\Delta_{R}$ not only sets the symmetry-breaking scale but also giving mass to the heavy $U(1)_{B-L}$ gauge boson $Z^{{}^{\prime}}$. In the split seesaw context, we show that the scale $\nu_{R}\simeq(10^{5}-3\times 10^{13})$ GeV that is consistent with both GW observations and evades the collider bounds [11]. The heavy mass of the gauge boson suppresses its emission from the string network, favoring decay channels such as GWs or sterile neutrino emission. This suppression is essential in our model, where we explore the cosmological signatures of such strings without conflicting with astrophysical constraints.
In this work, we analyze the formation and evolution of type-I CSs in the framework of the split seesaw mechanism embedded in an LR symmetric model. We discuss the implications for neutrino mass generation and CS phenomenology, and identify benchmark parameter sets consistent with current theoretical and observational bounds. Our results demonstrate the compatibility of this unified framework with high-energy particle physics and early universe cosmology. This regime is often overlooked in the literature, as most cosmic string studies focus on type-II strings, where parallel strings tend to repel at large distances while antiparallel strings attract [7].
The unique synergy of this framework is threefold: Firstly, the split seesaw provides a natural explanation for small neutrino masses and motivates the small quartic coupling necessary for type-I strings. Secondly, the left-right symmetry-breaking scale $\nu_{R}$ is simultaneously the source of heavy neutrino masses and the string tension. Thirdly, this connects the scale of neutrino mass generation directly to an observable cosmological signal-GWs from CSs—offering a novel, multi-messenger probe of neutrino physics. The leptogenesis in the LR model has been discussed in Ref. [12] and some analysis of GW production in the LR model has been discussed in Ref. [13, 14, 15, 16].
The paper is organized as follows: in Sec. II, we discuss the LR symmetric model and the formation of CSs as a result of symmetry breaking, we define the Lagrangian and scalar potential in Sec. II.1, in Sec. II.2 we discuss the neutrino mass generation in the split seesaw framework. We briefly discuss type-1 CSs and their phenomenological implications in Sec. III and lastly in Sec. IV we summarize our conclusions.
II Cosmic Strings in a Left-Right Symmetric Model
We consider a left-right (LR) symmetric extension of the Standard Model based on the gauge group:
$G_{\text{LR}}=SU(3)_{C}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L},$ (1)
which naturally incorporates the seesaw mechanism for neutrino mass generation. The field content of the LR model is defined in Tab. 1
Field Symbol Representation
LH quark doublet $Q_{L}$ $(\mathbf{3},\mathbf{2},\mathbf{1},+\tfrac{1}{3})$
RH quark doublet $Q_{R}$ $(\mathbf{3},\mathbf{1},\mathbf{2},+\tfrac{1}{3})$
LH lepton doublet $L_{L}$ $(\mathbf{1},\mathbf{2},\mathbf{1},-1)$
RH lepton doublet $L_{R}$ $(\mathbf{1},\mathbf{1},\mathbf{2},-1)$
Higgs bidoublet $\Phi$ $(\mathbf{1},\mathbf{2},\mathbf{2},0)$
Left Higgs triplet $\Delta_{L}$ $(\mathbf{1},\mathbf{3},\mathbf{1},+2)$
Right Higgs triplet $\Delta_{R}$ $(\mathbf{1},\mathbf{1},\mathbf{3},+2)$
Table 1: Field content of the minimal left-right symmetric model with representations under $SU(3)_{C}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L}$.
The symmetry is broken by a RH scalar triplet:
$\Delta_{R}\sim(1,1,3,+2),$ (2)
which acquires a VEV, spontaneously breaking the gauge group via:
$SU(2)_{R}\times U(1)_{B-L}\xrightarrow{\langle\Delta_{R}\rangle}U(1)_{Y}.$ (3)
This breaking leads to a non-trivial vacuum manifold
$\mathcal{M}=\frac{SU(2)_{R}\times U(1)_{B-L}}{U(1)_{Y}},$ (4)
with the first homotopy group
$\pi_{1}(\mathcal{M})=\mathbb{Z},$ (5)
indicating the formation of topologically stable CSs. For a detailed discussion on the formation of CSs, we refer the reader to [17, 18].
II.1 Lagrangian and Scalar Potential
The relevant kinetic and potential terms for $\Delta_{R}$ are given by:
$\mathcal{L}\supset\text{Tr}[(D_{\mu}\Delta_{R})^{\dagger}(D^{\mu}\Delta_{R})]-V(\Delta_{R}),$ (6)
with the covariant derivative defined as:
$D_{\mu}\Delta_{R}=\partial_{\mu}\Delta_{R}-ig_{R}[W_{R\mu},\Delta_{R}]-i2g_{B-L}B_{\mu}\Delta_{R},$ (7)
where $\partial_{\mu}$ is the ordinary derivative, $g_{R},\,W_{R_{\mu}}$ is the associated gauge coupling and bosons with $SU(2)_{R}$ respectively, $g_{B-L},\,B_{\mu}$ is the associated gauge coupling and bosons of $U(1)_{B-L}$. The scalar potential for $\Delta_{R}$ is,
$V(\Delta_{R})=-\mu^{2}\,\text{Tr}[\Delta_{R}^{\dagger}\Delta_{R}]+\lambda\left(\text{Tr}[\Delta_{R}^{\dagger}\Delta_{R}]\right)^{2}.$ (8)
Here $\mu$ is a positive mass-squared parameter that triggers spontaneous symmetry breaking when the potential is minimized, and $\lambda$ is the quartic self-coupling constant. This potential leads to a non-zero VEV for $\Delta_{R}$, breaking $SU(2)_{R}\times U(1)_{B-L}\rightarrow U(1)_{Y}$ and producing Majorana masses for RH neutrinos via the seesaw mechanism [19, 20, 21]. Assuming a vacuum configuration:
$\langle\Delta_{R}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0\\ v_{R}&0\end{pmatrix}$, (9)
which is a triplet under the gauge group $SU(2)_{R}$. This triplet is used to break the RH gauge symmetry $SU(2)_{R}\times U(1)_{B-L}$ down to the Standard Model’s hypercharge group $U(1)_{Y}$. This breaking gives rise to the formation of CSs. In the Abelian-Higgs model, the string tension is defined to be
$\mu\simeq 2\pi B(\beta)\nu_{R}^{2}.$ (10)
Here $\nu_{R}$ is the VEV of the scalar field $\Delta_{R}$, $\beta\equiv m_{\Delta_{R}}^{2}/m_{W_{R}}^{2}$ is the ratio between the mass square of the scalar and gauge field. The function $B(\beta)$ encodes the dependence on the scalar-to-gauge mass ratio. While its asymptotic behavior is well-known ($B(\beta)\rightarrow 1$ for $\beta\ll 1$ and $B(\beta)\sim ln(\beta)$ for $\beta\gg 1$) [22, 23], we adopt a phenomenological interpolation formula that accurately approximates numerical solutions across the full range of $\beta$ [24, 25]:
$B(\beta)\sim\dfrac{1}{1+2.5\left(\beta^{-1}-1\right)^{1.2}}.$ (11)
This fit is chosen for its simplicity and accuracy in representing the transition between the type-I ($\beta<1$) and type-II ($\beta>1$) regimes, which is crucial for our analysis.
II.2 Mass Generation in Split Seesaw
In split seesaw models, the smallness of the neutrino Yukawa coupling arises from localization in extra dimensions [26, 4]. In the split seesaw framework, we assume the existence of a single extra spatial dimension $y\in[0,L]$. The Standard Model fields are localized on a 4D brane at $y=0$, while the RH neutrinos $N_{i}$ propagate in the bulk. The 5D wavefunction of the $i$-th RH neutrino has the form:
$\Psi_{N_{i}}(x,y)=\psi_{N_{i}}(x)f_{i}(y),\quad\text{with}\quad f_{i}(y)=\sqrt{2M_{i}}e^{-M_{i}y},$ (12)
where $M_{i}$ is the 5D bulk mass term, assumed positive. The 5D Yukawa coupling to the Higgs field $\Phi$ localized at $y=0$ generates a suppressed effective 4D Yukawa coupling:
$Y^{\text{eff}}_{i\alpha}=Y^{(5D)}_{i\alpha}\cdot f_{i}(0)=Y^{(5D)}_{i\alpha}\cdot\sqrt{2M_{i}}e^{-M_{i}L}.$ (13)
The Majorana mass arises from a coupling of $\Psi_{N_{i}}$ to the $\Delta_{R}$ field on the brane at $y=L$:
$M_{N_{i}}\sim v_{R}.$ (14)
The Dirac mass term is:
$m_{D}=Y^{\text{eff}}v_{\text{EW}}\sim e^{-M_{i}L}v_{\text{EW}}.$ (15)
The standard type-I seesaw formula then gives:
$M_{\nu}\simeq-m_{D}M_{N}^{-1}m_{D}^{T}\sim\left(e^{-M_{i}L}\right)^{2}\cdot\frac{v_{\text{EW}}^{2}}{v_{R}},$ (16)
naturally yielding small active neutrino masses without requiring tiny dimensionless couplings.
The exponential suppression $e^{-M_{i}L}$ requires $L>1/M_{i}$ to be significant. However, to ensure the validity of the 44D effective field theory (EFT) at the temperature of the $SU(2)_{R}\times U(1)_{B-L}$ phase transition ($T\sim\nu_{R}$), we must impose that the compactification scale, $M_{c}\sim 1/L$ is significantly higher than $\nu_{R}$. This avoids Kaluza-Klein mode excitations and higher-dimensional operators disrupting the dynamics of CS formation. Consequently, we require
$\nu_{R}\ll M_{c}<M_{i}.$ (17)
This hierarchy ensures that the 44D description of the phase transition and the resulting topological defects remain robust. The small Yukawa couplings are generated by the large bulk mass terms $M_{i}$, not an excessively large extra dimension that would compromise the EFT. In Fig. 1, we have shown the neutrino mass evolution in the split seesaw mechanism versus the symmetry-breaking scale. The colored shaded region indicates the NANOGrav [27] excluded regime as illustrated in Fig. 2 via frequency bins. The inside colored curves indicate the different values of the exponential suppression factor. The colored stars correspond to the atmospheric scale mass, $M_{\nu}\simeq 0.05$ eV for each exponential suppression factor. The colliders constrain the lower value $\nu_{R}\gtrsim 10^{5}$ GeV [11].
Figure 1: Neutrino mass evolution versus symmetry-breaking scale has been presented. The colored shaded region indicates the NANOGrav [27] excluded regime as shown in Fig. 2. The colored curves represent the different values of the exponential suppression factor. The colored stars correspond to the atmospheric scale mass, $M_{\nu}\simeq 0.05$ eV for each exponential suppression factor. The colliders constrain the lower value $\nu_{R}\gtrsim 10^{5}$ GeV [11].
III Type-I Cosmic Strings
A key feature of our framework is the natural emergence of type-I CSs, characterized by $\beta=\lambda/(2g^{2})<1$. This is achieved through a small quartic coupling $\lambda$, which is well-motivated within the split seesaw mechanism. In conventional 44D seesaw models, the small neutrino Yukawa coupling often requires fine-tuning. In our higher-dimensional setup, the small effective Yukawa is generated geometrically by wavefunction suppression, Eq. 13. This liberates the quartic coupling $\lambda$ from the obligation to be large. In fact, a small value of $\lambda$ is technically natural. This naturalness is explicitly realized in supersymmetric (SUSY) extensions. In SUSY theories, the quartic couplings of scalar fields are not free parameters but are determined by the gauge couplings, leading to a natural expectation of $\lambda\sim O(0.1)$ [28]. While our present construction is non-SUSY, it is motivated by the fact that a small $\lambda$ is a robust feature of well-motivated UV completions. This distinguishes our scenario from a finely-tuned one and provides a strong theoretical motivation for exploring the $\beta<1$ regime.
Furthermore, the value of $\beta$ profoundly impacts the CS network evolution and its GW signature. While the expression for $\Omega_{\text{GW}}(f)$ in Eq. 18 has a generic form, it depends critically on the string tension $\mu\propto\nu_{R}^{2}\,B(\beta)$. The function $B(\beta)$ varies from $1$ for $\beta\ll 1$ to $ln(\beta)$ for $\beta\gg 1$, leading to a significantly different normalization of the GW spectrum for the same symmetry-breaking scale $\nu_{R}$. More importantly, the parameter $\beta$ determines the fraction of energy lost to gravitational radiation versus particle emission, which influences the loop number density (parameterized by $\zeta$ and $\alpha$ in Eq. 18) [29, 30]. Type-I strings have attractive self-interactions, leading to a different loop-chopping efficiency and network scaling solution compared to the more commonly studied type-II case. Our analysis, which focuses on the $\beta\ll 1$ regime, therefore explores a distinct and observationally viable corner of the CS parameter space that is intimately linked to the neutrino mass generation mechanism. Our analysis is now extended to the case $\beta\simeq 0.01$.
III.1 Phenomenological Implications
Let us now explain the GW power spectrum for type-I cosmic strings. We follow the treatment given in [10]. The total GW density at the present frequency $f$ is,
$\Omega_{\rm GW}(f)$ $\simeq\dfrac{2\,\Gamma\,G\mu^{2}}{\rho_{c}\,f}\,\dfrac{\zeta}{\alpha\left(\alpha+\Gamma\,G\mu\right)}\times\int_{t_{f}}^{t_{0}}d\tilde{t}\left(\dfrac{a(\tilde{t})}{a(t_{0})}\right)^{5}$ (18)
$\left(\dfrac{a(t_{i})}{a(\tilde{t})}\right)^{3}\,t_{i}^{-4}.$
Here $\Gamma\simeq 50$ quantifies the efficiency of GW emission, $G\mu$ is the dimensionless string tension, $\rho_{c}=3H_{0}^{2}/(8\pi G)$ is the critical density ($H_{0}$ is the Hubble parameter today and $G$ is Newton’s constant). We considered the largest width of the loop parameterized by $\alpha=0.1$ and $\zeta=10$, which quantifies the net energy flux associated with the loops. The upper limit on the integration is up to the present time $t_{0}$, the lower limit $t_{f}$ is the time when the cosmic strings enter the scaling regime, $a(t)$ being the scale factor and $t_{i}$ is the loop formation time given as,
$t_{i}=\dfrac{\tilde{l}+\Gamma\,G\mu\,\tilde{t}}{\alpha+\Gamma\,G\mu}.$ (19)
Here $\tilde{l}=2a(\tilde{t})/(fa(t_{0}))$ is the size of the loop. We have presented the GW power spectra in Fig. 2, which indicates the testability of the model and the symmetry-breaking scale in the present and future planned experiments. The different colors in the power spectrum indicate the benchmark point for the symmetry-breaking scale as indicated in Fig. 1. To be consistent with the current PTA dataset [27], the symmetry-breaking scale $\nu_{R}\lesssim 3\times 10^{13}$ GeV as shown by the gray dashed line in Fig. 2. Therefore, the emission of GWs from type-I CSs provides testable signatures of neutrino masses through cosmology.
Figure 2: The evolution of GW spectra for type-I CSs for different values of $\nu_{R}$ corresponding to the benchmark points given in Fig. 1 and indicated in the inset. The dashed gray line corresponds to $\nu_{R}\simeq 3\times 10^{13}$, consistent with the PTAs. The colored shaded regions indicate the sensitivity curves of present (solid boundaries) LIGO O3 [31], NANOGrav [27] and future (dashed boundaries) LIGO O5, SKA [32], THEIA [33], LISA [34], $\mu$-ARES [35], BBO [36], U-DECIGO [37, 38], CE [39] and ET [40] experiments.
IV Conclusions
This work adopts the left-right symmetric model that provides a natural gauge-invariant embedding of right-handed neutrinos as part of $SU(2)_{R}$ doublets. By incorporating the split seesaw mechanism into this framework, one achieves exponentially suppressed effective Yukawa couplings due to wavefunction localization in extra dimensions, while the heavy Majorana mass $M_{N}$ arises from the spontaneous breaking of $SU(2)_{R}\times U(1)_{B-L}$. This unified picture not only explains the smallness of neutrino masses without fine-tuning but also predicts the formation of type-I cosmic strings with $\beta\ll 1$ associated with the BLB–L breaking scale. In the split seesaw mechanism, one can explain the matter-antimatter asymmetry of the universe, as well as dark matter, which we leave for future work, but some discussion has already been made in Ref. [26]. This framework, therefore, simultaneously addresses neutrino physics and early-universe cosmology in a consistent high-energy theory.
Acknowledgments
The author gratefully acknowledges Qaisar Shafi, Rishav Roshan, Rinku Maji, and Anish Ghoshal for their valuable discussions.